# Universidad Autónoma de Madrid, Cantoblanco, E-28049 Madrid, Spain

### v1

@@ 1 Introduction
The TEK model is an effective framework to study lattice gauge theories in the large-$N$ limit by exploiting a fundamental property of large-$N$ gauge theories: the so-called large-$N$ volume reduction PhysRevLett.48.1063 ; BHANOT198247 ; Gross:1982at ; GONZALEZARROYO1983174 ; PhysRevD.27.2397 ; Aldazabal:1983ec ; Kiskis:2002gr ; Narayanan:2003fc ; Kovtun:2007py ; Unsal:2008ch ; Gonzalez-Arroyo:2010omx ; Neuberger:2020wpx . In their pioneering paper, Eguchi and Kawai PhysRevLett.48.1063 showed that, in the large-$N$ limit, the Yang–Mills lattice theory enjoys a dynamical equivalence between space-time and color degrees of freedom, leading the model to become insensitive to finite-volume effects when $N\to\infty$. Taking this idea to the extreme, it allows to study large-$N$ gauge theories in the thermodynamic limit even on a volume-reduced 1-site lattice Gonzalez-Arroyo:1983cyv ; Das:1984jh ; Das:1985xc ; Kiskis:2009rf ; Hietanen:2009ex ; Hietanen:2010fx ; Bringoltz:2011by ; Hietanen:2012ma ; Gonzalez-Arroyo:2012euf ; Gonzalez-Arroyo:2013bta ; Lohmayer:2013spa ; Gonzalez-Arroyo:2014dua ; GarciaPerez:2014azn ; GarciaPerez:2015rda ; Perez:2015ssa ; Gonzalez-Arroyo:2015bya ; Perez:2017jyq ; GarciaPerez:2020gnf ; Perez:2020vbn ; Butti:2023hfp ; Bonanno:2023ypf ; Butti:2022sgy ; Bonanno:2024bqg ; Bonanno:2024onr . This has the advantage of dramatically reducing the computational burden related to the space-time degrees of freedom, and allows to reach very large values of $N\sim\mathcal{O}(10^{2}-10^{3})$ in feasible simulations. Thus, unlike standard approaches to study large-$N$ gauge theories—where the $N=\infty$ limit is typically approached from $N\lesssim 10$ values via extrapolation DelDebbio:2001sj ; Lucini:2001ej ; DelDebbio:2002xa ; Lucini:2004my ; DelDebbio:2006yuf ; Vicari:2008jw ; Allton:2008ty ; Lucini:2010nv ; Lucini:2012gg ; Bali:2013kia ; Bonati:2016tvi ; Ce:2016awn ; Bennett:2020hqd ; Bonanno:2020hht ; Athenodorou:2021qvs ; Bennett:2022gdz ; Bonanno:2022yjr ; Bonanno:2024ggk ; Athenodorou:2024loq ; Sharifian:2025fyl ; DeGrand:2016pur ; Hernandez:2019qed ; DeGrand:2020utq ; Hernandez:2020tbc ; DeGrand:2021zjw ; Baeza-Ballesteros:2022azb ; DeGrand:2023hzz ; DeGrand:2024lvp ; DeGrand:2024frm ; Baeza-Ballesteros:2025iee ; Butti:2025rnu —the TEK model practically allows to work directly at $N=\infty$. Large-$N$ volume reduction works provided that center symmetry is unbroken. Spontaneous center symmetry breaking that would occur when reducing the lattice size due to the well-known deconfinement transition Fingberg:1992ju ; Beinlich:1997ia ; Campostrini:1998zd ; Lucini:2001ej ; Lucini:2002ku ; Lucini:2003zr ; Lucini:2004yh ; Lucini:2005vg ; Lucini:2012wq ; Borsanyi:2022xml ; Lucini:2023irm ; Cohen:2023hbq is avoided in the TEK approach by imposing twisted boundary conditions on the gauge fields in all directions, hence the name of the model. [^1] This is not the only strategy that has been applied to enforce center symmetry on a small box and achieve large-$N$ volume independence see, e.g., Refs. BHANOT198247 ; Narayanan:2003fc ; Kovtun:2007py ; Unsal:2008ch . The TEK approach has been extensively used in the last decade, and has allowed to achieve significant progress in the study of several large-$N$ gauge theories, such as large-$N$ QCD Gonzalez-Arroyo:2012euf ; GarciaPerez:2014azn ; Gonzalez-Arroyo:2015bya ; Perez:2020vbn ; Bonanno:2023ypf ; Butti:2023hfp ; Bonanno:2025hzr , large-$N$ SUSY Yang–Mills Butti:2022sgy ; Bonanno:2024bqg ; Bonanno:2024onr , and large-$N$ adjoint QCD Gonzalez-Arroyo:2013bta ; GarciaPerez:2015rda ; Hamada:2025whg .
@@ 2 Random Matrix Theory predictions
These two routes should of course lead to compatible estimates. Since all RMT predictions are described by the same single parameter $\Sigma$, we will check that different matching prescriptions—i.e., different choices of kk—to extract the chiral condensate give compatible result for the condensate when the corresponding $\lambda_{k}$ eigenvalues are in the RMT regime. Moreover, we will compare these determinations of $\Sigma$ with the one found in Bonanno:2025hzr for non-chiral Wilson quarks.
@@ 3.2 Lattice chiral Dirac operator
In order to restore physical units in the lattice operator $D$ and in the lattice mass $\hat{m}$ and connect them to their physical counterparts, one needs to perform the following rescaling:
$\not${D} $=$ $\frac{M(2-M)}{a}D_{0}(M)$ (32)
@@
where $m$ is the bare quark mass in physical units. The pre-factor M(2M)M(2-M), which stems from the choice of the kernel in Eq. (27), was worked out analytically by requiring that the spectrum of the massless operator in Eq. (32) in the free case was equal to $\not${p} in momentum space. More details can be found in the Appendix (see also Ref. Aoki:2001su ). Clearly, the same normalization applies when restoring physical units in the eigenvalues of the massless overlap Dirac operator:
@@ 4.3 Parameter-dependent RMT predictions
NN $b$ $\tilde{\chi}^{2}$ $p$-value
$10^{3}\times a^{3}\Sigma/N$
[[RMT fit]
$10^{3}\times a^{3}\Sigma/N$
[[Eq. (47)]
@@
$b$ NN
$10^{3}\times a^{3}\Sigma/N$
[…]
$10^{3}\times a^{3}\Sigma/N$
[…]
$10^{3}\times a^{3}\Sigma/N$
[…]
$10^{3}\times a^{3}\Sigma/N$
@@
The differences observed at smaller values of $N$ in the condensates extracted from higher eigenvalues, as discussed in the previous section, can be regarded as finite-volume effects due to not being yet in the regime where RMT provides a reliable description. It is known that RMT is expected to hold in the so-called $\epsilon$-regime of QCD Gasser:1987ah , see, e.g., Refs. Verbaarschot:2000dy ; Basile:2007ki . Customarily, QCD is studied in the so-called $p$-regime, obtained taking first the thermodynamic limit $L\to\infty$ at fixed and finite quark mass and then the chiral limit, so that $m_{\pi}L\to\infty$. In the $\varepsilon$-regime, instead, $m\to 0$ and $L\to\infty$ with $m\Sigma V\lesssim\mathcal{O}(1)$. Since $m\Sigma V=2(F_{\pi}m_{\pi}L^{2})^{2}$, this means that $m_{\pi}L\to 0$ while $F_{\pi}L\to\infty$. It was shown in Ref. Gasser:1987ah that finite-volume corrections to the thermodynamic limit in the $\epsilon$-regime are power-like, and can be expanded in series of the variable $1/(F_{\pi}^{2}L^{2})=1/(F_{\pi}^{2}\sqrt{V})$. This behavior is very different from the customary exponential suppression of finite-size effects in the $p$-regime, and is related to the peculiar way of taking the chiral and thermodynamic limits.
@@ 4.4 Renormalization of the overlap condensate and comparison with Wilson quarks
This will be our reference matching point to compute $Z_{\scriptscriptstyle{\rm S}}$ from Eq. (55), and is reported in Fig. 6 as a solid line. This value is obtained from the customary exponential best fit to the pion correlator, cf. Fig. 7 (left panel). This correlation function is obtained from a standard GEVP analysis using an extended basis of smeared pseudo-scalar operators. As a standard cross-check, we verified that the effective mass, obtained by solving the GEVP at all times, exhibits a plateau in the same range where we performed the exponential best fit. A constant fit to this plateau gives the perfectly compatible result $\frac{m_{\pi}^{\rm(ref)}}{\sqrt{\sigma}}=0.98(4)$, cf. Fig. 7 (right panel). For more details on the GEVP analysis we performed, we refer the reader to Ref. Bonanno:2025hzr , a dedicated paper on the extraction of the low-lying meson spectrum of large-$N$ QCD from the TEK model.
@@ Acknowledgements
It is a pleasure to thank Carlos Pena for many fruitful discussions and for reading this manuscript. We also acknowledge useful discussions with Pilar Hernández and Rajamani Narayanan. This work is partially supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grant PID2021-127526NB-I00, both funded by MCIN/AEI/10.13039/ 501100011033. It is also partially funded by the European Commission – NextGenerationEU, through Momentum CSIC Programme: Develop Your Digital Talent. K.-I. I. is supported in part by MEXT as “Feasibility studies for the next-generation computing infrastructure”. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We acknowledge support from the SFT Scientific Initiative of INFN. This work was partially supported by the Simons Foundation grant 994300 (Simons Collaboration on Confinement and QCD Strings). Numerical calculations have been performed on the Finisterrae III cluster at CESGA (Centro de Supercomputación de Galicia), on the Drago cluster at CSIC (Consejo Superior de Investigaciones Científicas) and on the Hydra and Ciclope clusters at IFT. We acknowledge HPC support by Emilio Ambite, staff hired under the Generation D initiative, promoted by Red.es, an organization attached to the Ministry for Digital Transformation and the Civil Service, for the attraction and retention of talent through grants and training contracts, financed by the Recovery, Transformation and Resilience Plan through the European Union’s Next Generation funds. We have also used computational resource of Oakbridge-CX, at the University of Tokyo through the HPCI System Research Project (Project ID: hp230021 and hp220011), of Cygnus at Center for Computational Sciences, University of Tsukuba, of SQUID at D3 Center of Osaka university through the RCNP joint use program.
@@ Appendix A The spectrum of the free overlap operator
where $s_{\mu}=\sin(ap_{\mu})$, $s^{2}=\sum_{\mu}\sin^{2}(ap_{\mu})$ and $\hat{s}^{2}=\sum_{\mu}\sin^{2}(ap_{\mu}/2)$, with $p_{\mu}=2\pi n_{\mu}/\sqrt{N}$, $n_{\mu}=0,\cdots,\sqrt{N}-1$. These results agree with those quoted for the standard overlap operator with identical kernel in Ref. Aoki:2001su .

### v2

@@ 1 Introduction
The TEK model is an effective framework to study lattice gauge theories in the large-$N$ limit by exploiting a fundamental property of large-$N$ gauge theories: the so-called large-$N$ volume reduction PhysRevLett.48.1063 ; BHANOT198247 ; Gross:1982at ; GONZALEZARROYO1983174 ; PhysRevD.27.2397 ; Aldazabal:1983ec ; Kiskis:2002gr ; Narayanan:2003fc ; Kovtun:2007py ; Unsal:2008ch ; Gonzalez-Arroyo:2010omx ; Neuberger:2020wpx . In their pioneering paper, Eguchi and Kawai PhysRevLett.48.1063 showed that, in the large-$N$ limit, the Yang–Mills lattice theory enjoys a dynamical equivalence between space-time and color degrees of freedom, leading the model to become insensitive to finite-volume effects when $N\to\infty$. Taking this idea to the extreme, it allows to study large-$N$ gauge theories in the thermodynamic limit even on a volume-reduced 1-site lattice Gonzalez-Arroyo:1983cyv ; Das:1984jh ; Das:1985xc ; Kiskis:2009rf ; Hietanen:2009ex ; Hietanen:2010fx ; Bringoltz:2011by ; Hietanen:2012ma ; Gonzalez-Arroyo:2012euf ; Gonzalez-Arroyo:2013bta ; Lohmayer:2013spa ; Gonzalez-Arroyo:2014dua ; GarciaPerez:2014azn ; GarciaPerez:2015rda ; Perez:2015ssa ; Gonzalez-Arroyo:2015bya ; Perez:2017jyq ; GarciaPerez:2020gnf ; Perez:2020vbn ; Butti:2023hfp ; Bonanno:2023ypf ; Butti:2022sgy ; Bonanno:2024bqg ; Bonanno:2024onr . This has the advantage of dramatically reducing the computational burden related to the space-time degrees of freedom, and allows to reach very large values of $N\sim\mathcal{O}(10^{2}-10^{3})$ in feasible simulations. Thus, unlike standard approaches to study large-$N$ gauge theories — where the $N=\infty$ limit is typically approached from $N\lesssim 10$ values via extrapolation DelDebbio:2001sj ; Lucini:2001ej ; DelDebbio:2002xa ; Lucini:2004my ; DelDebbio:2006yuf ; Vicari:2008jw ; Allton:2008ty ; Lucini:2010nv ; Lucini:2012gg ; Bali:2013kia ; Bonati:2016tvi ; Ce:2016awn ; Bennett:2020hqd ; Bonanno:2020hht ; Athenodorou:2021qvs ; Bennett:2022gdz ; Bonanno:2022yjr ; Bonanno:2024ggk ; Athenodorou:2024loq ; Sharifian:2025fyl ; DeGrand:2016pur ; Hernandez:2019qed ; DeGrand:2020utq ; Hernandez:2020tbc ; DeGrand:2021zjw ; Baeza-Ballesteros:2022azb ; DeGrand:2023hzz ; DeGrand:2024lvp ; DeGrand:2024frm ; Baeza-Ballesteros:2025iee ; Butti:2025rnu — the TEK model practically allows to work directly at $N=\infty$. Large-$N$ volume reduction works provided that center symmetry is unbroken. Spontaneous center symmetry breaking that would occur when reducing the lattice size due to the well-known deconfinement transition Fingberg:1992ju ; Beinlich:1997ia ; Campostrini:1998zd ; Lucini:2001ej ; Lucini:2002ku ; Lucini:2003zr ; Lucini:2004yh ; Lucini:2005vg ; Lucini:2012wq ; Borsanyi:2022xml ; Lucini:2023irm ; Cohen:2023hbq is avoided in the TEK approach by imposing twisted boundary conditions on the gauge fields in all directions, hence the name of the model. [^1] This is not the only strategy that has been applied to enforce center symmetry on a small box and achieve large-$N$ volume independence see, e.g., Refs. BHANOT198247 ; Narayanan:2003fc ; Kovtun:2007py ; Unsal:2008ch .
In order to clarify how our method compares with more traditional ones, it is useful to recall that the nature of finite-$N$ corrections is different in the TEK model than in standard approaches. The model at finite $N$ is actually related to the corresponding non-commutative field theory models Gonzalez-Arroyo:1983zev ; Connes:1987ue ; Rieffel:1998vs ; Ambjorn:2000cs ; Douglas:2001ba ; Alvarez-Gaume:2001czc . The form of these corrections is already clear from the results in the original paper PhysRevD.27.2397 . The planar diagram contribution appears as the one formulated in a finite lattice of size $(\sqrt{N})^{4}$. Thus, as it will be done in this paper, the size of the lattice in physical units has to be taken large enough to make these corrections small for the observable in question. This, however, is not different from the minimal volume requirements in more conventional approaches. Indeed, one can take advantage of the latter studies to quantify the importance of these corrections. In addition, finite-$N$ values in the TEK model also affect the suppression of non-planar contributions. This effect has been studied extensively in more recent times starting in Ref. Gonzalez-Arroyo:2010omx . It turns out that the choice of the flux parameter $k$, appearing in the twist factor implementing twisted boundary conditions in the TEK action (see Sec. 3), plays a crucial role in keeping these corrections small. An appropriate $N$-dependent choice $k(N)$ gives rise to corrections of order $1/N^{2}$ with a small prefactor. This, given the large values of $N$ achievable in the model, makes these corrections negligibly small. All these effects have been studied in detail in the recent literature both perturbatively Perez:2017jyq ; Gonzalez-Arroyo:2019zfm and non-perturbatively Gonzalez-Arroyo:2014dua . As a matter of fact, the TEK approach has been extensively used in lattice simulations in the last decade, and has allowed to achieve significant progress in the study of several large-$N$ gauge theories, such as large-$N$ QCD Gonzalez-Arroyo:2012euf ; GarciaPerez:2014azn ; Gonzalez-Arroyo:2015bya ; Perez:2020vbn ; Bonanno:2023ypf ; Butti:2023hfp ; Bonanno:2025hzr , large-$N$ SUSY Yang–Mills Butti:2022sgy ; Bonanno:2024bqg ; Bonanno:2024onr , and large-$N$ adjoint QCD Gonzalez-Arroyo:2013bta ; GarciaPerez:2015rda ; Hamada:2025whg .
@@ 2 Random Matrix Theory predictions
These two routes should of course lead to compatible estimates. Since all RMT predictions are described by the same single parameter $\Sigma$, we will check that different matching prescriptions — i.e., different choices of $k$ — to extract the chiral condensate give compatible result for the condensate when the corresponding $\lambda_{k}$ eigenvalues are in the RMT regime. Moreover, we will compare these determinations of $\Sigma$ with the one found in Bonanno:2025hzr for non-chiral Wilson quarks.
@@ 3.2 Lattice chiral Dirac operator
In order to restore physical units in the lattice operator $D$ and in the lattice mass $\hat{m}$ and connect them to their physical counterparts, one needs to perform the following rescalings:
$\not${D} $=$ $\frac{M(2-M)}{a}D_{0}(M),$ (32)
@@
where $m$ is the bare quark mass in physical units. The prefactor M(2M)M(2-M), which stems from the choice of the kernel in Eq. (27), was worked out analytically by requiring that the spectrum of the massless operator in Eq. (32) in the free case was equal to $\not${p} in momentum space. More details can be found in the Appendix (see also Ref. Aoki:2001su ). Clearly, the same normalization applies when restoring physical units in the eigenvalues of the massless overlap Dirac operator:
@@ 4.3 Parameter-dependent RMT predictions
NN $b$ $\tilde{\chi}^{2}$ $p$-value $10^{3}\times a^{3}\Sigma/N$ [[RMT fit] $10^{3}\times a^{3}\Sigma/N$ [[Eq. (47)]
@@
The differences observed at smaller values of $N$ in the condensates extracted from higher eigenvalues, as discussed in the previous section, can be regarded as finite-volume effects due to not being yet in the regime where RMT provides a reliable description. It is known that RMT is expected to hold in the so-called $\epsilon$-regime of QCD Gasser:1987ah , see, e.g., Refs. Verbaarschot:2000dy ; Basile:2007ki . Customarily, QCD is studied in the so-called $p$-regime, obtained taking first the thermodynamic limit $L\to\infty$ at fixed and finite quark mass and then the chiral limit, so that $m_{\pi}L\to\infty$. In the $\varepsilon$-regime, instead, $m\to 0$ and $L\to\infty$ at fixed $m\Sigma V\lesssim\mathcal{O}(1)$. Since $m\Sigma V=2(F_{\pi}m_{\pi}L^{2})^{2}$, this means that $m_{\pi}L\to 0$ while $F_{\pi}L\to\infty$. It was shown in Ref. Gasser:1987ah that finite-volume corrections to the thermodynamic limit in the $\epsilon$-regime are power-like, and can be expanded in series of the variable $1/(F_{\pi}^{2}L^{2})=1/(F_{\pi}^{2}\sqrt{V})$. This behavior is very different from the customary exponential suppression of finite-size effects in the $p$-regime, and is related to the peculiar way of taking the chiral and thermodynamic limits.
@@ 4.4 Renormalization of the overlap condensate and comparison with Wilson quarks
This will be our reference matching point to compute $Z_{\scriptscriptstyle{\rm S}}$ from Eq. (55), and is reported in Fig. 6 as a solid line. This value is obtained from the customary exponential best fit to the pion correlator, cf. Fig. 7 (left panel). This correlation function is obtained from a standard GEVP analysis using an extended basis of smeared pseudo-scalar operators. As a standard cross-check, we verified that the effective mass, obtained by solving the GEVP at all times, exhibits a plateau in the same range where we performed the exponential best fit. A constant fit to this plateau gives the perfectly compatible result $\frac{m_{\pi}^{\rm(ref)}}{\sqrt{\sigma}}=0.98(4)$, cf. Fig. 7 (right panel). For more details on the GEVP analysis we performed, we refer the reader to Ref. Bonanno:2025hzr , a dedicated paper on the extraction of the low-lying meson spectrum of large-$N$ QCD from the TEK model. Note also that, since in Ref. Bonanno:2025hzr we did not observe any finite-volume effect in meson masses obtained for $m_{\pi}\ell\gtrsim 4$ within our percent precision, we expect that any possible $1/V$ finite-volume effect due to the fixed $Q=0$ topology Brower:2003yx ; Aoki:2007ka of our gauge configurations is below our statistical accuracy.
@@ Acknowledgements
It is a pleasure to thank Carlos Pena for many fruitful discussions and for reading this manuscript. We also acknowledge useful discussions with Pilar Hernández and Rajamani Narayanan. This work is partially supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by the grants No. PID2021-127526NB-I00 and PID2024-160152NB-I00, both funded by MCIN/AEI/10.13039/ 501100011033. It is also partially funded by the European Commission – NextGenerationEU, through Momentum CSIC Programme: Develop Your Digital Talent. K.-I. I. is supported in part by MEXT as “Feasibility studies for the next-generation computing infrastructure”. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We acknowledge support from the SFT Scientific Initiative of INFN. This work was partially supported by the Simons Foundation grant 994300 (Simons Collaboration on Confinement and QCD Strings). Numerical calculations have been performed on the Finisterrae III cluster at CESGA (Centro de Supercomputación de Galicia), on the Drago cluster at CSIC (Consejo Superior de Investigaciones Científicas) and on the Hydra and Ciclope clusters at IFT. We acknowledge HPC support by Emilio Ambite, staff hired under the Generation D initiative, promoted by Red.es, an organization attached to the Ministry for Digital Transformation and the Civil Service, for the attraction and retention of talent through grants and training contracts, financed by the Recovery, Transformation and Resilience Plan through the European Union’s Next Generation funds. We have also used computational resource of Oakbridge-CX, at the University of Tokyo through the HPCI System Research Project (Project ID: hp230021 and hp220011), of Cygnus at Center for Computational Sciences, University of Tsukuba, of SQUID at D3 Center of Osaka university through the RCNP joint use program.
@@ Appendix A The spectrum of the free overlap operator
where $s_{\mu}=\sin(ap_{\mu})$, $s^{2}=\sum_{\mu}\sin^{2}(ap_{\mu})$ and $\hat{s}^{2}=\sum_{\mu}\sin^{2}(ap_{\mu}/2)$, with $ap_{\mu}=2\pi n_{\mu}/\sqrt{N}$, $n_{\mu}=0,\cdots,\sqrt{N}-1$. These results agree with those quoted for the standard overlap operator with identical kernel in Ref. Aoki:2001su .